OpenStudy (anonymous):

The seats at a local baseball stadium are arranged so that each row has five more seats than the row in front of it. If there are four seats in the first row, how many total seats are in the first 24 rows? @kc_kennylau

OpenStudy (solomonzelman):

\[a_{n}=a_{n-1}+5\]

OpenStudy (solomonzelman):

So the difference is 5.

OpenStudy (solomonzelman):

A1 = 4

OpenStudy (solomonzelman):

Do you know how to find a sum of an arithmetic series?

OpenStudy (anonymous):

Where did you get the formula from @SolomonZelman

OpenStudy (solomonzelman):

Which one the first thing I wrote?

OpenStudy (anonymous):

Yes that

OpenStudy (solomonzelman):

Well, because each next row is 5 seats more than the previous row.

OpenStudy (solomonzelman):

The 24th row would be \[A_n=a_1+d(n-1)\]in your case, for the 24th row, \[A_{24}=4+5(24-1)\]

OpenStudy (anonymous):

207?

OpenStudy (solomonzelman):

No, 119. After you find the 224th row, use this: \[S_n=\frac{n(a_1+a_n)}{2}\]

OpenStudy (solomonzelman):

I meant the 24th not 224th.

OpenStudy (anonymous):

BUT 9*23=204

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

why 9 ?

OpenStudy (anonymous):

5+4=9

OpenStudy (solomonzelman):

\[S_n=\frac{n(a_1+a_2)}{2}\]\[S_{24}=\frac{24(4+119)}{2}\]

OpenStudy (solomonzelman):

4+5=9 would be just the second row, we don;t need to know that.

OpenStudy (anonymous):

1476? @SolomonZelman

OpenStudy (solomonzelman):

Yup!

OpenStudy (anonymous):

Thank you!!! @SolomonZelman

OpenStudy (solomonzelman):

Anytime, !